please help me solve equation

4a over 3 +5 = a over 2

Thanks from Edgar Delgodo

Answer by stanbon(75887)

4a over 3 +5 = a over 2
--------------------------------------
(4a/3) + 5 = (a/2)
---
Multiply thru by 6 to get:
8a + 30 = 3a
5a = -30
a = -6
==============
Cheers,
Stan H.